OpenStudy (anonymous):

Wirte an equation for the line that passes through (0,0) and (2,3).

OpenStudy (anonymous):

(y2-y1)/(x2-x1) will get you your slope: (3-0)/(2-0) = 3/2

OpenStudy (anonymous):

using the equation y - y1 = slope(x - x1) you get y - 3 = (3/2)(x - 2)

OpenStudy (anonymous):

you might be more comfortable with the y = mx +b form: y = (3/2)x

OpenStudy (anonymous):

so, is the equation y=(3/2)x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how would you get that answer by using the y=mx+b format?

OpenStudy (anonymous):

well, you know that m = (3/2), so then you can plug in either (2,3) or (0,0) in for y and x and then solve for b. Using (0,0) is easiest because you find that b = 0 with little effort

OpenStudy (anonymous):

or the y2-y1/x2-x1 format

OpenStudy (anonymous):

y2-y1/x2-x1 is used to find your slope, so it would be (3-0)/(2-0) = 3/2

OpenStudy (anonymous):

Wirte an equation for the line that passes through (1,0) and (3,4).

OpenStudy (anonymous):

Then what is the equation for finding the b?

OpenStudy (anonymous):

b = y-mx

OpenStudy (anonymous):

thanks